OpenStudy (anonymous):

Solve system of equations: 4x+12y=5 -x-3y=-2

OpenStudy (ranga):

Multiply the second equation by 4 and add it to the first. That will get rid of x and you can solve for y. Put the y value in the second equation and solve for x.

OpenStudy (anonymous):

When I plug the second equation -4x-12y=-8

OpenStudy (anonymous):

is that part right

OpenStudy (ranga):

Yes that part is correct.

OpenStudy (anonymous):

cool

OpenStudy (ranga):

Okay, this is a trick question because the -4x - 12y = -8 that you got by multiplying the second equation by 4 can be rewritten as: 4x + 12y = 8 (by multiplying by -1) Compare it to the first equation that says: 4x+12y = 5 How can the same 4x + 2y be 8 and also be 5. That means there is no solution for this set of equations.

OpenStudy (anonymous):

How did you come up with multiplying by -1? There has to be a answer because this a homework question.

OpenStudy (austinl):

Multiply the second equation by 4, add the top and bottom equations together. What do you get?

OpenStudy (ranga):

Okay. Let me start from the beginning. 4x+12y=5 -x-3y=-2 Multiply equation (2) by 4: -4x - 12y = -8 Add it to the first equation: 4x+12y=5 -4x - 12y = -8 Add them: 0 + 0 = -3 Whenever you get answers like this it means there is NO value of x and y that can satisfy both equations at the same time. Therefore, there is no solution for this problem. You can answer: There is no solution for x and y and your teacher will say that is the correct answer.

OpenStudy (anonymous):

Austin L I got -4x-12y=-8

OpenStudy (austinl):

now add that to the top equation.

OpenStudy (anonymous):

ok